Help! how do I entered this using the BA II Plus Texas Instrument calculator? Jeffery Wei received a 7-year non-subsidized student loan of $32,000 at an annual interest rate of 5.6%. What are Jeffery's monthly loan payments for this loan after he graduates in 4 years? (Round your answer to the nearest cent.)
Answer #1

ANSWER:

I = 5.6% PER YEAR OR 5.6% / 12 PER MONTH

N = 4 YEARS OR 48 MONTHS

MONTHLY LOAN PAYMENTS = LOAN AMOUNT(A/P,I,N) = 32,000(A/P,5.6%/12,48) = 32,000 * 0.0233 = 745.67

SO THE MONTHLY LOAN PAYMENTS ARE $745.67 OR $745.70 (ROUNDED OFF TO 1 DECIMAL PLACE.)